# Fairness for the People, by the People: Minority Collective Action

### v1

@@ 5 Limitations of Minority Collective Action
Previous work on collective action assumes that the collective is uniformly sampled from the distribution $\mathbb{P}_{0}$ and that the collective has a perfect oracle for the conditional distribution $\mathbb{P}_{0}{}\left(Y|X\right)$. Yet, our method restricts collective participation to minority members and approximates this conditional distribution. Those differences introduce limitations to the existing theory, which we analyze and theoretically quantify in this section.
@@
As mentioned above, we focus on collectives composed solely of minority members, unlike prior work. This restriction expresses scenarios in which majority members lack incentives to support changes that would benefit the minority, and instead prefer to preserve the status quo. Naturally, this limitation reduces the collective’s impact, as demonstrated in the following example.
Consider a binary classification task on the two-dimensional 4-Gaussian mixture model $\mathbb{P}_{\text{4GMM}}$ where each Gaussian belongs to a distinct combination of label and group membership, as illustrated in Figure˜6. Each label consists of a large majority subgroup and a significantly smaller minority subgroup. We can then state the following informal result about the EqOd fairness violation of ERM.
@@
Consider a dataset sampled from the distribution $\mathbb{P}_{\text{4GMM}}$ described above, where every minority point participates in the collective action by flipping all $y=0$ labels to $y=1$. Then, under sufficiently separable clusters, with high probability, the EqOd of the ERM classifier minimizing the logistic loss will asymptotically approach $0.5$.
The formal Proposition˜5 is provided in Section˜B.2 along with all necessary assumptions, which holds for a broader family of distributions and can be extended to any dimensionality $\mathbb{R}^{d}$ using techniques similar to those in Chaudhuri et al. [35]. Although Proposition˜2 is not a formal lower bound, it emphasizes an important limitation: collective action restricted to the minority cannot generally achieve perfect fairness, even under very advantageous conditions involving a maximum-sized collective, a strong strategy, and a disregard for accuracy. This limitation stands in contrast to standard firm-side bias mitigation methods, which in principle, achieve perfect fairness.
There are several reasons why relabeling alone may not be enough to get perfect fairness. For one, relabeling according to the counterfactual implicitly assumes that the label is determined by the same features across the majority and minority, but this assumption is not valid under certain distribution shift between the groups. To illustrate, consider a firm training a classifier to screen candidates for a managerial position. Majority members may be more educated, while minority members may have more hands-on experience rather than formal education. In this case, a counterfactual label associated with the majority is disjointed from the features associated with the features, rendering the signal erasure strategy irrelevant. Future work could study this problem and determine when it is beneficial to change the features as well.
We empirically corroborate the findings of Proposition˜2 on real world datasets by examining the fairness–accuracy tradeoff of several fair learning methods. Figure˜7 compares the Pareto fronts of RB-prob, one of our minority collective action methods, with established firm-side methods. We observe that the lowest fairness violation achievable by RB-prob is greater than that of the firm-side approaches. However, the firm-side methods are able to arrive at perfect fairness only at a cost of prohibitively high prediction error. But, inspecting the region where the error is small compared to the base classifier, the fairness of RB-prob is comparable to that of the firm-side methods.
@@
Figure 7: User-side method cannot achieve perfect fairness, while the firm-side pre-processing method FARE [24] and the post-processing method calibrated equalized odds [36] attain 0 EqOd with large error. However, RB-prob’s fairness is better than the base classifier, with a smaller error than the firm-side methods.
@@
$\rho:=\mathbb{P}_{0}\left(\mathcal{A}\left(x\right)\neq\arg\max_{y^{\prime}}\mathbb{P}_{0}\left[y^{\prime}|g\left(x\right)\right]\right).$ (10)
Given this definition, we derive the following lower bound on success, proved in Section˜B.3.
@@
Figure 8: The Pareto fronts for using a fair representation when computing the KNN for RB-dist dominate the Pareto fronts for KNN computed on untransformed features. The blue stars represent the KNN without transforming the data, and the yellow triangles represent the KNN when the data is transformed using FARE [24]. The lines are fitted by a polynomial of degree 2 to guide the eye.
Impact of feature representations
Since the methods RB-label and RB-dist rely on KNN, their performance is sensitive to the choice of distance metric and feature representation. In our main experiments, we used Euclidean distance in the original feature space, which is convenient but could be suboptimal. Here, we explore whether FRL can learn a more suitable representation space for KNN. A fair representation maps the data into a space where the group-based bias is removed while preserving informative features. Intuitively, such representations may help RB-label and RB-dist to better estimate the counterfactual labels.
To formalize this intuition, we consider predicting the counterfactual label of minority points using a $1$-NN classifier on majority data, i.e., assigning each minority point the label of its nearest neighbor in the majority. In settings where the minority is distributed differently than the majority (e.g., $\mathbb{P}_{\text{4GMM}}$), this task can be challenging. The following informal result compares the error of $1$-NN in the original features space to its error in a learned fair representation.
Proposition 4 (Informal).
The formal statement, Theorem˜1, with the proof and assumptions can be found in Section˜B.4. The result suggests that FRL can reduce the counterfactual label error $\rho$ of RB-label and RB-dist, consequently improving the lower bound of the collective’s success according to Proposition˜3. Empirically, Figure˜8 indicates that applying FARE [24] before the KNN step improves the Pareto front for RB-dist. On the other hand, methods that rely purely on predictive information, such as RB-prob, can perform worse, due to FRL inadvertently removing features predictive of the class label.
@@ 6 Related work
Optimizing for fairness metrics often comes at the cost of reduced classification accuracy, leading to the well-documented accuracy–fairness tradeoff [2, 3, 4, 5]. In response, previous work has proposed fairness interventions at different stages of the ML pipeline: pre-processing methods modify the training data before learning [6, 7, 8, 24], in-processing methods adjust the learning algorithm itself [37, 38, 33, 39], and post-processing methods correct the predictions of a trained (unfair) classifier [19, 40, 41, 42]. A firm can introduce any of these categories into its pipeline, while users, who control only their data can only partially implement pre-processing methods. However, as mentioned in Section˜2.2, using feature-changing pre-processing methods such fair representation learning [8, 24] demand changing those features during inference time as well.
Still, a couple of pre-processing methods suggest changing only the labels, similarly to our proposed collective action. The method by Luong et al. [7] compares between the minority KNN and majority KNN and flip the labels according to the difference of positive labels between the two groups of neighbors. This method resembles RB-label, with the difference that RB-label examines only the majority KNN in order to approximate the counterfactual. Similarly, the approach of Kamiran and Calders [6] trains a regressor to predict $y=1$ outcome probabilities, and flip the label of minority members with $y=0$ labels and high probability according to the regressor to have $y=1$, and similarly flip majority $y=1$ labels to $y=0$. Flipping from both groups is done to preserve the error of the classifier. Our method RB-prob differs by training the regressor only on the majority to better approximate the counterfactuals. Since this approach requires flipping the labels of majority members as well, it cannot be completey adopted by the collective. In Section˜D.1 we compare the between RB-prob to CND and KDP, and find that our method, based on the counterfactual search, is more efficient in terms of the required number of label flips.
@@ 7 Conclusion
This work demonstrates that user-side methods, specifically minority collective action, can effectively reduce unfairness in machine learning. While much of the existing fairness research focused on firm-side methods, paradoxically these often come at a cost that may not be worth to the firm. This catch emphasizes the importance of studying user-side approaches for bias mitigation. We also note that in general, collective action methods can be exploited by malicious parties seeking self-gain or harming other communities, and it is important to be discussing these limitations and possibly regulate them.
We introduce three practical methods that a collective can easily implement to relabel itself, and show empirically that collective action can considerably reduce unfairness in a variety of datasets, though not completely. Importantly, we also examine the limitations of a minority being composed of only minority members, and how the success is affected by approximating the counterfactual labels.
Our proposed methods require the collective to change their own labels, an action that often comes with a price, as the users have to go against their true nature. However, as has been studied on real world cases, minority members willingly participate in collective action to benefit their demographic, after being encouraged by their community [16] or by the media [15].
Overall, this paper shows a practical use case of collective action in the hopes of sparking further research into applications of collective action and user-side methods for social good.
@@ Acknowledgments
Appendix A Preliminaries
A.1 Statistical parity and equalized odds
@@
where a smaller SP value indicates fairer treatment across groups. However, SP does not account for the ground-truth labels $y$, and thus optimizing for SP can degrade the overall accuracy. For example, a classifier that always predicts $\hat{y}=1$ will have perfect SP but a high prediction error. Alternatively, a stricter notion called equalized odds (EqOd) [19] requires that both the true positive rate and false positive rate be equal across groups. Here the EqOd difference is defined as
@@
A.2 Suboptimal Bayes classifier
@@
$\mathrm{TV}\left(\cdot,\cdot\right)$ is the total variation distance between two distributions. The definition is discussed more in Hardt et al. [10].
Appendix B Theoretical Results and Proofs
B.1 Counterfactual fairness as success
@@
B.2 Impossibility of fairness under ERM
The following proposition follows the structure of Theorem 6 in Chaudhuri et al. [35]. For a vector $x\in\mathbb{R}^{d}$, let $D\left(x\right)$ denote a distribution on $\mathbb{R}^{d}$ with mean $x$. Let $p$ and $m$ be the number of majority and minority sample, respectively with $p\gg m$.
Assumption 1 (Concentration Condition, Assumption 2 from Chaudhuri et al. [35]).
@@ B.2 Impossibility of fairness under ERM
Let $B,A$ be two sets of points that are sampled from $D(0)$. We will always associate $B$ with negatively labelled points and $A$ with positive, as will be clear below. Following Chaudhuri et al. [35], define the sets
@@
As Chaudhuri et al. [35] shows, the ERM solution can be written as $w^{*}=\alpha^{*}\hat{\mu}+\sigma\beta^{*}\hat{\psi}$, where
@@
$F(\alpha)=\min_{\alpha}\Big{\{}$ $\max\big{(}f_{1}\left(\alpha\right)-\alpha\|\mu\|+\sigma\beta\|\psi\|,\ f_{2,-}\left(\alpha\right)-\alpha\|\mu\|-\sigma\beta\|\psi\|,\ f_{4,-}\left(\alpha\right)-\alpha\|\mu\|-\sigma\beta\|\psi\|\big{)}$
$+\max\big{(}f_{3}\left(\alpha\right)-\alpha\|\mu\|+\sigma\beta\|\psi\|,\ f_{2,+}\left(\alpha\right)-\alpha\|\mu\|-\sigma\beta\|\psi\|,\ f_{4,+}\left(\alpha\right)-\alpha\|\mu\|-\sigma\beta\|\psi\|\big{)}\Big{\}}.$
@@
$f_{1}(\alpha)-f_{2,\pm}(\alpha)\geq 2\|\psi\|,\quad f_{1}(\alpha)-f_{4,\pm}(\alpha)\geq 2\|\psi\|,\quad f_{3}(\alpha)-f_{2,\pm}(\alpha)\geq 2\|\psi\|,\quad f_{3}(\alpha)-f_{4,\pm}(\alpha)\geq 2\|\psi\|.$ (30)
@@
$\big{[}f_{1}(\alpha)-\alpha\|\mu\|+\beta\|\psi\|\big{]}-\big{[}f_{2,-}(\alpha)-\alpha\|\mu\|-\beta\|\psi\|\big{]}=\big{(}f_{1}(\alpha)-f_{2,-}(\alpha)\big{)}+2\beta\|\psi\|\ \geq\ 2\|\psi\|+2\beta\|\psi\|\ >\ 0,$
@@
$\big{[}f_{1}(\alpha)-\alpha\|\mu\|-\beta\|\psi\|\big{]}-\big{[}f_{2,-}(\alpha)-\alpha\|\mu\|+\beta\|\psi\|\big{]}=\big{(}f_{1}(\alpha)-f_{2,-}(\alpha)\big{)}-2\beta\|\psi\|\ \geq\ 2\|\psi\|-2\beta\|\psi\|\ \geq 0,$
@@
Lemma 1 (Approximate Maximization Lemma - I, Lemma 14 from Chaudhuri et al. [35]).
@@
This result can also be extended to $\mathbb{R}^{d}$ using techniques similar to those in Chaudhuri et al. [35]. This result also encompasses the 4-Gaussian mixture model $\mathbb{P}_{\text{4GMM}}$ used in Section˜5 as a special case, leading to the following. See 2
B.3 Success Bound With Label Error
The following proof uses Lemma 11 from Hardt et al. [10].
Lemma 2 (Lemma 11 from Hardt et al. [10]).

### v2

@@ 5 Limitations of Minority Collective Action
Previous work on ACA assumes that the collective is uniformly sampled from the distribution $\mathbb{P}_{0}$ and that the collective has a perfect oracle for the conditional distribution $\mathbb{P}_{0}{}\left(Y|X\right)$. Yet, our method restricts collective participation to minority members and approximates this conditional distribution. Those differences introduce limitations to the existing theory, which we analyze and theoretically quantify in this section.
@@
As mentioned above, we focus on collectives composed solely of minority members, unlike prior work. This restriction expresses scenarios in which majority members lack incentives to support changes that would benefit the minority, and instead prefer to preserve the status quo. Naturally, this limitation reduces the collective’s impact.Consider a binary classification task on the two-dimensional 4-Gaussian mixture model $\mathbb{P}_{\text{4GMM}}$ where each Gaussian belongs to a distinct combination of label and group membership, as illustrated in Figure˜6. Each label consists of a large majority subgroup and a significantly smaller minority subgroup. We can then state the following informal result about the EqOd fairness violation of ERM.
@@
Consider the dataset $\mathbb{P}_{\text{4GMM}}$, where every minority point participates in the ACA by flipping all $y=0$ labels to $y=1$. Then, under sufficiently separable clusters, with high probability, the EqOd of the ERM classifier minimizing the logistic loss will asymptotically approach $0.5$.
The formal Proposition˜5 is provided in Section˜C.2 along with all necessary assumptions, which holds for a broader family of distributions and can be extended to any dimensionality $\mathbb{R}^{d}$ using techniques similar to those in chaudhuriWhyDoesThrowing2023. Although Proposition˜2 is not a formal lower bound, it emphasizes an important limitation: ACA restricted to the minority cannot generally achieve perfect fairness, even under very advantageous conditions involving a maximum-sized collective, a strong strategy, and a disregard for accuracy. This limitation stands in contrast to standard firm-side bias mitigation methods, which in principle, achieve perfect fairness. There are several reasons why relabeling alone may not be enough to get perfect fairness. For one, relabeling according to the counterfactual implicitly assumes that the label is determined by the same features across the majority and minority, but this assumption is not valid under certain distribution shift between the groups. To illustrate, consider a firm training a classifier to screen candidates for a managerial position. Majority members may be more educated, while minority members may have more hands-on experience rather than formal education. In this case, a counterfactual label associated with the majority is disjointed from the features associated with the features, rendering the signal erasure strategy irrelevant. Future work could study this problem and determine when it is beneficial to change the features as well.
We empirically corroborate the findings of Proposition˜2 on real world datasets by examining the fairness–accuracy tradeoff of several fair learning methods. Figure˜7 compares the Pareto fronts of RB-prob, one of our minority ACA methods, with established firm-side methods. We observe that the lowest fairness violation achievable by RB-prob is greater than that of the firm-side approaches. However, the firm-side methods are able to arrive at perfect fairness only at a cost of prohibitively high prediction error. But, inspecting the region where the error is small compared to the base classifier, the fairness of RB-prob is comparable to that of the firm-side methods.
@@
Figure 7: User-side method cannot achieve perfect fairness, while the firm-side pre-processing method FARE [jovanovicFAREProvablyFair2023] and the post-processing method calibrated equalized odds [pleissFairnessCalibration2017] attain 0 EqOd with large error. However, RB-prob’s fairness is better than the base classifier, with a smaller error than the firm-side methods.
@@
$\rho:=\mathbb{P}_{0}(\mathcal{A}\left(x\right)\neq\arg\max_{y^{\prime}}\mathbb{P}_{0}\left[y^{\prime}|g\left(x\right)\right]).$ (10)
Given this definition, we derive the following lower bound on success, proved in Section˜C.3.
@@
Impact of feature representations
Since the methods RB-label and RB-dist rely on KNN, their performance is sensitive to the choice of distance metric and feature representation. In our main experiments, we used Euclidean distance in the original feature space, which is convenient but could be suboptimal. Here, we explore whether FRL can learn a more suitable representation space for KNN. A fair representation maps the data into a space where the group-based bias is removed while preserving informative features. Intuitively, such representations may help RB-label and RB-dist to better estimate the counterfactual labels. To formalize this intuition, we consider predicting the counterfactual label of minority points using a $1$-NN classifier on majority data, i.e., assigning each minority point the label of its nearest neighbor in the majority. In settings where the minority is distributed differently than the majority (e.g., $\mathbb{P}_{\text{4GMM}}$), this task can be challenging. The following informal result compares the error of $1$-NN in the original features space to its error in FRL.
Proposition 4 (Informal).
The formal statement, Theorem˜1 can be found in Section˜C.4. The result suggests that FRL can reduce the counterfactual label error $\rho$ of RB-label and RB-dist, consequently improving the lower bound of the collective’s success according to Proposition˜3. Empirically, Figure˜8 indicates that applying FARE [jovanovicFAREProvablyFair2023] before the KNN step improves the Pareto front for RB-dist. On the other hand, methods that rely purely on predictive information, such as RB-prob, can perform worse, due to FRL inadvertently removing features predictive of the class label.
@@
Figure 8: The Pareto fronts for using a fair representation when computing the KNN for RB-dist dominate the Pareto fronts for KNN computed on untransformed features. The blue stars represent the KNN without transforming the data, and the yellow triangles represent the KNN when the data is transformed using FARE [jovanovicFAREProvablyFair2023]. The lines are fitted to guide the eye.
@@ 6 Related work
Optimizing for fairness often comes at the cost of reduced classification accuracy, leading to the well-documented accuracy–fairness tradeoff [menonCostFairnessBinary2018, zhaoInherentTradeoffsLearning2019, dehdashtianUtilityfairnessTradeoffsHow2024, sadeghiCharacterizingTradeoffInvariant2022]. In response, previous work has proposed fairness interventions at different stages of the ML pipeline: pre-processing methods modify the training data before learning [kamiranClassifyingDiscriminating2009, luongKNNImplementationSituation2011, zemelLearningFairRepresentations2013, jovanovicFAREProvablyFair2023], in-processing methods adjust the learning algorithm itself [agarwalReductionsApproachFair2018, namLearningFailureDebiasing2020, sagawaDistributionallyRobustNeural2020, liuJustTrainTwice2021], and post-processing methods correct the predictions of a trained (unfair) classifier [hardtEqualityOpportunitySupervised2016, alghamdiAdultCOMPASFair2022, tifreaFRAPPEGroupFairness2024, cruzUnprocessingSevenYears2024]. A firm can introduce any of these categories into its pipeline, while users, who control only their data can only partially implement pre-processing methods. However, as mentioned in Section˜2.2, using feature-changing pre-processing methods such fair representation learning [zemelLearningFairRepresentations2013, jovanovicFAREProvablyFair2023] demand changing those features during inference time as well.
Still, some pre-processing methods change only the labels, similarly to our proposed collective action. The method by luongKNNImplementationSituation2011 compares between the minority KNN and majority KNN and flip the labels according to the difference of positive labels between the two groups of neighbors. This method resembles RB-label, with the difference that RB-label examines only the majority KNN in order to approximate the counterfactual. The approach of kamiranClassifyingDiscriminating2009 trains a regressor to predict $y=1$ outcome probabilities, and flip the label of minority members with $y=0$ labels and high probability according to the regressor to have $y=1$, and similarly flip majority $y=1$ labels to $y=0$. Flipping from both groups supposedly preserves the error of the classifier. Our method RB-prob differs by training the regressor only on the majority to better approximate the counterfactuals. Since this approach requires flipping the labels of majority members as well, it cannot be completey adopted by the collective. In Section˜E.1 we compare the between RB-prob to CND and KDP, and find that our method is more efficient in terms of number of label flips.
@@ 7 Conclusion
This work demonstrates that user-side methods, can effectively reduce unfairness in machine learning. While much of the existing fairness research focused on firm-side methods, these often come at a cost that may not be worth to the firm. This catch emphasizes the importance of studying user-side approaches for bias mitigation. We introduce three practical methods that a collective can easily implement to relabel itself, and show empirically that ACA can considerably reduce unfairness in a variety of datasets, though not completely. Importantly, we also examine the limitations of a minority being composed of only minority members, and how the success is affected by approximating the counterfactual labels. Our proposed methods require the collective to change their own labels, which often comes with a price, as the users have to go against their true nature. However, as has been studied on real world cases, minority members willingly participate in collective action to benefit their demographic, after being encouraged by their community [begenyPowerIngroupPromoting2022] or by the media [saleemWhenHowNegative2021].
We also note that in general, ACA methods can be exploited by malicious parties seeking self-gain or harming other communities, and it is important to discuss these limitations and possibly regulate them. Overall, this paper shows a practical use case of collective action in the hopes of sparking further research into applications of ACA and user-side methods for social good.
@@ Acknowledgments
8 Appendix
Appendix A Motivating examples
To recognize real-world problems that lend themselves well to collective action for fairness one needs to look for the following few characteristics:
  • Firm and goal: A firm trains a predictive model primarily to minimize average error, with little incentive to protect minority groups.
  • End-users: People who use the platform and whose behavior generates data for the firm’s dataset.
  • Disadvantaged group: A subgroup of end-users who is treated unfairly.
  • Relabeling possibility: How the minority can relabel themselves to make the trained classifier fairer.
Here we suggest several cases that answer these characteristics.
  1. 1.
    Content moderation
    • Firm and goal: A global social-media company optimizes a high-recall harmful-content detector measured on its largest user pools.
    • End-users: Everyday users of the platform who can flag offensive content.
    • Disadvantaged group: Slurs, insults, or cultural references specific to minority communities are not flagged often enough, so the model fails to detects harmful content in those groups’ languages.
    • Relabeling possibility: The minority flags borderline content from their community that the platform’s global guidelines ignore.
  2. 2.
    Resume screening
    • Firm and goal: A multi-national HR firm trains a classifier to extract skills from resumes.
    • End-users: Job applicants submitting resumes.
    • Disadvantaged group: Applicants from a disadvantaged minority may lack formal education and degrees compared to the majority, but may have informal training which the classifier ignores.
    • Relabeling possibility: Applicants can reframe their work experience, e.g. framing working at a store as being a salesperson, or managing shifts as managerial experience.
  3. 3.
    Medical treatment prediction
    • Firm and goal: A nationwide insurer builds a treatment-recommendation model to minimize average costs and adverse events.
    • End-users: Patients who report their treatment outcomes (pain levels, recovery time, side effects).
    • Disadvantaged group: Minority groups may experience different side effects or recovery rates than the majority, so the model recommends suboptimal treatments for them.
    • Relabeling possibility: Individual patients record more detailed outcomes rather than underreporting, e.g., consistently marking “still in pain” instead of “fine”.
  4. 4.
    Credit scoring
    • Firm and goal: A lender trains a credit-risk model to predict defaults and set loan terms, using historical repayment data.
    • End-users: Borrowers whose repayment or default becomes training labels.
    • Disadvantaged group: Disadvantaged groups may not have credit cards or have never taken loans, and only deal with cash but still pay their bills. These actions are “credit-invisible”.
    • Relabeling possibility: A borrower can report their payed bills, such as rent or utilities, as repaid loans. These become additional positive repayment labels.
  5. 5.
    Recommender systems
    • Firm and goal: A streaming platform trains recommender system to maximize engagement, heavily weighted toward mainstream content [baumannAlgorithmicCollectiveAction2024].
    • End-users: Users who like, skip, or re-listen to songs.
    • Disadvantaged group: Niche genres or local musicians get suppressed, as engagement data mostly comes from the majority’s preferences.
    • Relabeling possibility: Users can promote underrepresented content by repeatedly listening, liking, or playlisting it.
Appendix B Preliminaries
B.1 Statistical parity and equalized odds
@@
where a smaller SP value indicates fairer treatment across groups. However, SP does not account for the ground-truth labels $y$, and thus optimizing for SP can degrade the overall accuracy. For example, a classifier that always predicts $\hat{y}=1$ will have perfect SP but a high prediction error. Alternatively, a stricter notion called equalized odds (EqOd) [hardtEqualityOpportunitySupervised2016] requires that both the true positive rate and false positive rate be equal across groups. Here the EqOd difference is defined as
@@
B.2 Suboptimal Bayes classifier
@@
$\mathrm{TV}\left(\cdot,\cdot\right)$ is the total variation distance between two distributions. The definition is discussed more in hardtAlgorithmicCollectiveAction2023.
Appendix C Theoretical Results and Proofs
C.1 Counterfactual fairness as success
@@
C.2 Impossibility of fairness under ERM
The following proposition follows the structure of Theorem 6 in chaudhuriWhyDoesThrowing2023. For a vector $x\in\mathbb{R}^{d}$, let $D\left(x\right)$ denote a distribution on $\mathbb{R}^{d}$ with mean $x$. Let $p$ and $m$ be the number of majority and minority sample, respectively with $p\gg m$.
Assumption 1 (Concentration Condition, Assumption 2 from chaudhuriWhyDoesThrowing2023).
@@ C.2 Impossibility of fairness under ERM
Let $B,A$ be two sets of points that are sampled from $D(0)$. We will always associate $B$ with negatively labelled points and $A$ with positive, as will be clear below. Following chaudhuriWhyDoesThrowing2023, define the sets
@@
As chaudhuriWhyDoesThrowing2023 shows, the ERM solution can be written as $w^{*}=\alpha^{*}\hat{\mu}+\sigma\beta^{*}\hat{\psi}$, where
@@
$F(\alpha)=\min_{\alpha}\Big\{$ $\max\big(f_{1}\left(\alpha\right)-\alpha\|\mu\|+\sigma\beta\|\psi\|,$
$\phantom{\max\big(}f\left(\alpha\right)-\alpha\|\mu\|-\sigma\beta\|\psi\|,$
$\phantom{\max\big(}f_{4,-}\left(\alpha\right)-\alpha\|\mu\|-\sigma\beta\|\psi\|\big)$
$+\max\big(f_{3}\left(\alpha\right)-\alpha\|\mu\|+\sigma\beta\|\psi\|,$
$\phantom{+\max\big(}f_{2,+}\left(\alpha\right)-\alpha\|\mu\|-\sigma\beta\|\psi\|,$
$\phantom{+\max\big(}f_{4,+}\left(\alpha\right)-\alpha\|\mu\|-\sigma\beta\|\psi\|\big)\Big\}.$
@@
$f_{1}\left(\alpha\right)-f_{2,\pm}\left(\alpha\right)\geq 2\left\|\psi\right\|,$ $f_{1}\left(\alpha\right)-f_{4,\pm}\left(\alpha\right)\geq 2\left\|\psi\right\|,$ (30)
$f_{3}\left(\alpha\right)-f_{2,\pm}\left(\alpha\right)\geq 2\left\|\psi\right\|,$ $f_{3}\left(\alpha\right)-f_{4,\pm}\left(\alpha\right)\geq 2\left\|\psi\right\|.$
@@
$\left[f_{1}\left(\alpha\right)-\alpha\left\|\mu\right\|+\beta\left\|\psi\right\|\right]-\left[f_{2,-}\left(\alpha\right)-\alpha\left\|\mu\right\|-\beta\left\|\psi\right\|\right]$ $=\left(f_{1}\left(\alpha\right)-f_{2,-}\left(\alpha\right)\right)+2\beta\left\|\psi\right\|$
$\geq 2\left\|\psi\right\|+2\beta\left\|\psi\right\|$
$\geq 0$
@@
$\left[f_{1}\left(\alpha\right)-\alpha\left\|\mu\right\|-\beta\left\|\psi\right\|\right]-\left[f_{2,-}\left(\alpha\right)-\alpha\left\|\mu\right\|+\beta\left\|\psi\right\|\right]$ $=\left(f_{1}\left(\alpha\right)-f_{2,-}\left(\alpha\right)\right)-2\beta\left\|\psi\right\|$
$\geq 2\left\|\psi\right\|-2\beta\left\|\psi\right\|$
$\geq 0$
@@
Lemma 1 (Approximate Maximization Lemma - I, Lemma 14 from chaudhuriWhyDoesThrowing2023).
@@
This result can also be extended to $\mathbb{R}^{d}$ using techniques similar to those in chaudhuriWhyDoesThrowing2023. This result also encompasses the 4-Gaussian mixture model $\mathbb{P}_{\text{4GMM}}$ used in Section˜5 as a special case, leading to the following. See 2
C.3 Success Bound With Label Error
The following proof uses Lemma 11 from hardtAlgorithmicCollectiveAction2023.
Lemma 2 (Lemma 11 from hardtAlgorithmicCollectiveAction2023).